### v1

# VersiCode: Towards Version-controllable Code Generation

Abstract
Significant research has focused on improving the performance of large language model on code-related tasks due to their practical importance. Although performance is typically evaluated using public benchmark datasets, the existing datasets do not account for the concept of version, which is crucial in professional software development. In this paper, we introduce VersiCode, the first comprehensive dataset designed to assess the ability of large language models to generate verifiable code for specific library versions. VersiCode encompasses 300 libraries across more than 2,000 versions spanning 9 years. We design two dedicated evaluation tasks: version-specific code completion (VSCC) and version-aware code editing (VACE). Comprehensive experiments are conducted to benchmark the performance of LLMs, revealing the challenging nature of these tasks and VersiCode, that even state-of-the-art LLMs struggle to generate version-correct code. This dataset, together with the proposed tasks, sheds light on LLMs’ capabilities and limitations in handling version-specific code generation, and opens up an important new area of research for further investigation. The resources can be found at https://github.com/wutong8023/VersiCode.
1 Introduction
Large language models (LLMs) [35] and their code-specific variants, such as CodeLlaMa [37], have shown strong capabilities in code generation tasks. Many code generation datasets have been proposed [6, 27] to evaluate these models’ performance, which sees steady increases. While these datasets serve as important benchmarks, they all treat code generation as a static task [18], without considering the influence of versioning of dependent libraries. In contrast, in professional software development, as shown in Figure 1, dependency management is an essential practice that plays a key role in building dependable software. Keeping track of versions is a central step in dependency management. Given the evolution of software libraries and their APIs, two major challenges remain under-explored. 1) How to generate code that uses and fits a specific version of a library. 2) How to adapt code using an older version of a library to a newer version.
Figure 1: Two motivating scenarios for version-controllable code generation: (left) Interacting with LLMs in a browser, where slight query changes lead to incorrect answers, and (right) Programming in an IDE, explicitly specifying the version of dependency libraries.
In this paper, we propose two dedicated version-controllable code generation tasks to simulate realistic settings in professional software development: version-specific code completion (VSCC) and version-aware code editing (VACE), in which the versions of the libraries serve as a crucial constraint for code generation.
Existing code generation datasets, due to their obliviousness to version information, cannot support these two tasks [6, 27]. Thus, to evaluate LLMs’ performance on these two novel tasks, we curate and release VersiCode, a large-scale benchmark dataset in Python with comprehensive versioning metadata. VersiCode contains 300 libraries and 2,207 versions, spanning 9 years from 2015 to 2023. We carefully developed a human-LLM hybrid data collection and annotation pipeline to create a large and high-quality dataset. Each data instance in VersiCode is represented as a tuple of “<library version, functionality description, code snippets>”. Subsequently, by detecting version-sensitive parts (e.g. function or variable names), we post-processed the metadata into tasks of varying granularities: token-level, line-level, and block-level, reflecting different levels of difficulty by generating more content and reducing provided contextual cues. To the best of our knowledge, VersiCode is the first version-controllable benchmark dataset. Our new benchmark dataset VersiCode, viewed through the lens of library version evolution, addresses the following research questions:
  1. 1.
    How well do existing LLMs understand version requirements for code generation?
  2. 2.
    Can LLMs recall the version span of APIs in each library, i.e., from addition to deprecation? Furthermore, can LLMs reason across intermediate versions?
  3. 3.
    What is the trend of LLM performance regarding the release time of each library?
In our comprehensive experiments, we evaluated the performance of a large number of state-of-the-art LLMs, including open LLMs such as Llama 2 and Llama 3, proprietary LLMs such as GPT-3.5 and GPT-4o, as well as seven code-specific variants such as CodeLlama-13B, StarCoder2-15B, and CodeQwen1.5-7B. On the simplest token-level code completion task, these LLMs perform significantly worse on VersiCode than on other well-known datasets including HumanEval(+) and MBPP(+). For some LLMs, the performance gap of Pass@1 between existing benchmarks and VersiCode is more than 50 points. Even the most capable LLM, GPT-4o, obtains a Pass@1 score of 70.44 on VersiCode while achieving a score higher than 85 on HumanEval and MBPP. LLMs performance on line-level and block-level completion is in generally much lower than that on token-level. For the version-aware code editing task, there is significant room for improvement even for the most capable LLM GPT-4o. Our contributions are summarized as follows:
  • We propose two novel and important yet under-explored tasks in code generation, namely version-specific code completion (VSCC) and version-aware code editing (VACE).
  • We curate VersiCode, a new, large-scale version-controllable code generation dataset, aimed at comprehensively evaluating LLMs’ performance on the above two tasks.
  • We conduct comprehensive experiments on VersiCode, demonstrating that it is a high-quality and challenging dataset. The benchmarking results reveal that most LLMs struggle with version-specific code generation, especially with the latest libraries, highlighting areas for future exploration.
2 VersiCode: A Benchmark for Version-controllable Code Generation
VersiCode is a large-scale code generation benchmark dataset focusing on evolving library dependencies. We propose two tasks to simulate real-world applications: version-specific code completion and version-aware code editing, incorporating version information into code generation constraints. First, we discuss data curation, preprocessing of noisy code snippets and FAQs into organized metadata (Section 2.1). Based on this metadata, we describe the task design and quality control process (Section 2.2). We then address tagging API lifespan features per library version (Section 2.3). Finally, we provide data statistics for VersiCode and discuss future dataset extensions (Section 2.4).
2.1 Dataset Curation and Collection
As shown in Figure 2, we first collected permissively licensed Python repositories from GitHub that serve as the source code for Python libraries. These repositories are ranked by their popularity (as indicated by their collected stars). Using the list of popular libraries, we gathered data from three sources for each library: (1) Library Source Code: We collected all available versions of the library source code from GitHub, verifying with PyPI to ensure that the collected versions are formally released and can be installed via pip. From the library source code, we extracted official usage examples for each API from the docstrings. (2) Downstream Application Code: Given Python’s popularity in scientific programming, we collected the source code from top-tier research papers over 10 years as downstream applications. These applications are valuable due to being lightweight yet self-consistent, diverse in their topics, and tagged release timelines associated with publishing venues. Given the time span, this data source implicitly includes evolving libraries. (3) Stack Overflow: Using the library names as queries, we collected FAQ data from Stack Overflow, which provides real user queries and diverse user answers. We filtered the data to include only those queries that explicitly mention the versions of the libraries used, using heuristic rules, refer to Appendix A.
Given the high diversity and varied quality of the collected raw data, we adopted a hybrid annotation approach involving both human experts and LLMs, such as ChatGPT. (1) Library Source Code: The library version is concrete and explicitly available, but example usage varies across libraries and versions. We used an LLM with in-context learning to help extract example code from docstrings, preparing the library version and code snippets. (2) Downstream Applications: The version can easily be extracted from configuration files, typically named “requirements.txt”. We carefully filtered out Python files that are too long, do not mention the library version, or fail to compile. (3) Stack Overflow: Given the diversity of the questions, we designed strict heuristic rules to preliminarily annotate the library name, version, and corresponding Python code snippets mentioned in answers. We then distributed the pre-annotated data to six qualified human experts for verification and correction, ensuring the library version and code snippets are ready as well. With all pairs of library versions and code snippets, we employed ChatGPT with in-context learning to generate descriptions of the functionality for each code snippet. Each pair is wrapped in well-organized metadata.
Figure 2: The preprocessing pipeline to obtain metadata, structured as n-gram tuple of $<$library name, version, functionality description, code snippet$>$.
2.2 Task Design for Version-controllable Code Generation
As shown in Figure 3, we define each meta-instance $m_{i}=[l_{i};v_{i};d_{i};c_{i}]\in\mathcal{M}$, where $l$, $v$, $d$, and $c$ represent the library name, version, functionality description, and code snippet, respectively. We then design the following two version-controllable code generation tasks:
Version-Specific Code Completion (VSCC): Given a meta-instance $m_{i}$, the input is $x=[l_{i};v_{i};d_{i};c^{\prime}_{i}]$, where $c^{\prime}_{i}$ is the code snippet $c_{i}$ with selective masking, replacing the library- and version-sensitive contents with a special token. Depending on the length of the masked contents, the special token is defined as “[token-mask]”, “[line-mask]”, or “[block-mask]”, reflecting code completion on different granularity levels. The output $y$ is the masked content, typically containing function names or variables.
Version-Aware Code Editing (VACE): Given a pair of meta-instances $(m_{i},m_{j}|l_{i}==l_{j},d_{i}==d_{j},v_{i}\neq v_{j})$, the input $x=[l_{i};v_{i};d_{i};c_{i};v_{j}]$, and the output $y=c_{j}$. Note that version editing may require refactoring of the code structure, making it difficult to format as detailed as in token-level or line-level completion. Additionally, depending on the numerical relationship between $v_{i}$ and $v_{j}$, various scenarios arise, such as editing from an old version to a new version, or vice versa.
2.3 Lifespan Tagging of APIs
Consider an API $a$ added to the library $L$ in version $V_{s}$ and deprecated in version $V_{e}$, and is active in the intermediate version $V_{m}$ where $s\leq m\leq e$. We refer to the interval $[s,e)$ as the lifespan of $a$. To analyze model performance in detail, we assessed how up-to-date each LLM was concerning newly added or deprecated APIs per version. We compared the source code between any two consecutive versions of each library to detect changes in API or method names. Based on the detection results, we labeled the datasets obtained from the library source code as follows: “addition” indicates an API newly added in the current version and still applicable in subsequent versions; “deprecation” indicates the current version is the last usable version for the API; and “general” indicates the API usage method is inherited from the previous version.
Figure 3: The post-processing pipeline transforms metadata into specific tasks and the running example per task: (left) Leveraging pairs of metadata that share the same functionality but different library versions to construct block-level code editing instances; (right) Utilizing each metadata sample, masking version-sensitive content to create multi-granularity code completion instances.
2.4 Dataset Statistics, Scope, and Future Extensions
Table 1: Comparison between VersiCode and other code editing datasets, with VersiCode standing out as the largest annotated dataset specifically tailored for version adaptation.
Benchmark Source Language Samples Editing Task Granularity Collection Time Annotation
Defects4J [20] Open Source Programs Java 357 Debug Block-Level 2014 None
QuixBugs [26] Quixey Challenges Python, Java 40 Debug Line-Level 2017 Human
CoST [55] GeeksForGeeks Multi(=7) 132,046 Code Translation Line-Level, Block-Level 2022 None
XLCoST [54] GeeksForGeeks Multi(=8) 1,083,000 Code Translation Line-Level, Block-Level 2022 None
InstructCoder [15] Github Python 114,000 Code Refinement Block-Level 2023 LLM
MultilingualTrans [47] Programming Sites Multi(=8) 30,419 Code Translation Block-Level 2023 None
NicheTrans [47] Programming Sites Multi(>8) 236,468 Code Translation Block-Level 2023 None
LLMTrans [47] Hand-Written Multi(=8) 350 Code Translation Block-Level 2023 Human
Avatar [2] Programming Sites Python, Java 62,520 Code Translation Block-Level 2023 None
G-TransEval [19] Existing benchmark, GeeksForGeeks Multi(=5) 400 Code Translation Token-Level, Block-Level 2023 Human
EvalGPTFix [51] AtCoder Java 151 Debug Block-Level 2023 Human
DebugBench [41] LeetCode Multi(=3) 4,253 Debug Block-Level 2024 LLM
VersiCode Github Python 98,692 Version Adaptation Block-Level 2024 LLM
Dataset Statistics: We present the statistics of VersiCode in Table 2, using the NLTK  [^1] https://www.nltk.org/ tokenizer to compute the number of tokens. As shown in Table 1, we also provide a comparison between VersiCode and other code editing datasets. See Appendix B.2 for the comparison between VersiCode and other code completion datasets.
Scope: VersiCode supports version-specific code completion at the token, line, and block levels, enabling developers to navigate through version variations effortlessly. It also facilitates block-level version-aware code editing, empowering users to make precise modifications tailored to requirements of each version. The collected metadata also serves as a valuable resource for potential customized task modifications, aiding in fine-tuning workflows and enhancing model training for optimal performance.
Table 2: Data statistics of VersiCode.
# Num. of Libraries 300
# Num. of Versions 2,207
# Size of Metadata 11,268
# Task Type Completion Editing (old to new) Editing (new to old)
# Granularity Token Line Block Block Block
# Avg. Input Token 1,233 1,213 44 115 116
# Avg. Output Token 1 9 77 70 69
# Num. of Instances 13,532 13,530 1,618 49,346 49,346
Future Extension: VersiCode primarily supports Python, but we have also made efforts to extend it to multiple languages. Refer to the Appendix B.1 for comprehensive data statistics and initial multi-language evaluation results. Given the rapid evolution of software technologies, continued effort to further automate the curation pipeline and maintenance of VersiCode is needed.
3 Experiments
We undertake experiments to investigate the following research questions:
  1. 1.
    What is the quality and difficulty of VersiCode, particularly in comparison to related datasets and among diverse data sources?
  2. 2.
    What is the current proficiency of codeLLMs in generating version-controllable code, considering comparisons among multiple-granularity completions, for both code completion and code editing?
  3. 3.
    What is the timeliness of LLMs’ programming knowledge in code generation, with a focus on model performance categorized by year over the past decade?
3.1 Implementation Details
Models: We benchmarked VersiCode against popular open-domain LLMs and dedicated code-LLMs, including variant families such as GPT [35, 34, 36], LLaMa [42], Mistral [17], CodeLLaMa [37], CodeQwen [5], CodeGemma [7], StarCoder [28], Deepseek-Coder [13], and WizardCoder [32]. For smaller open-source models (e.g., < 20B parameters), we downloaded them from HuggingFace  [^2] https://huggingface.co/models and deployed them locally for inference. For larger models, such as LLaMa3 70B [33] and GPT-4o [36], we used their online APIs  [^3] https://together.ai  [^4] https://openai.com/ for inference.
Baseline Dataset: To assess the difficulty of VersiCode, we compared it with two well-known code generation datasets, HumanEval [27] and MBPP [18], and observed the overall performance of models. HumanEval [27] measures functional correctness in synthesizing programs from docstrings with 164 original problems, resembling simple software interview questions. MBPP [4], with about 1,000 crowd-sourced Python problems for entry-level programmers, covers programming fundamentals and standard library functionality, including task descriptions, code solutions, and three automated test cases for each problem. We also collected the evaluation results for their upgraded versions HumanEval+ [27] and MBPP+ [27]. Please refer to Appendix E.1 for details.
Data Preparation: As introduced in Section 2.2, we designed two types of version-controllable code generation tasks: version-specific code completion and version-aware code editing. The task granularities are categorized into token-level, line-level, and block-level to control difficulty and simulate different application scenarios. To better understand model performance, each instance in VersiCode is also tagged with the following: (1) data source, which includes library source code, downstream applications, and Stack Overflow (see Section 2.1); (2) feature type, including addition, deprecation, and general (see Section 2.3); (3) release time, i.e. the timestamp from GitHub and Stack Overflow); and (4) considering code editing instances constructed from pairs of metadata, the differences between source and target code versions result in various situations, such as updates from an older version to a newer version or vice versa. Additionally, we categorized versions according to version patterns, for example, treating torch v1.0.0 as a major version and torch v1.3.1 as a minor version, to identify combinations of major and minor version editing cases. These tags allow us to filter the evaluation dataset and gain sharper insights into model performance.
Evaluation Metrics: We use the following evaluation metrics for each task granularity: (1) Pass@$k$ for token-level generation [6]: For this metric, we generate $n\geq k$ samples per instance (with $n=100$ and $k\in\{1,3,10\}$ for our experiments). We count the number of correct samples $c\leq n$ judged by exact matching. Pass@$k$ is defined as the average performance over the task, calculated as $\mathbb{E}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right]$. (2) Identifier Sequence Match (ISM@$n$) and Prefix Match (PM@$n$) for line-level generation [1]: These metrics measure how closely the generated sequences match the ground truth. For block-level generation, we adopt the average performance over lines. Following the setup in Agrawal et al. [1], we generate $n=6$ independent samples per instance.
3.2 Results and Analysis
Figure 4: Pass@1 results for token-level code completion from VersiCode: (a1) comparison with existing benchmark datasets, (a2) performance grouped by data sources, and (b) performance grouped by API lifespan features.
Even token-level code completion is challenging. We present the pass@1 results of token-level code completion for LLMs on VersiCode, sorted by release time (see Figure 4-a1, highlighted in green). When compared to the Pass@1 results on HumanEval (in blue) and MBPP (in orange), all models perform significantly worse on VersiCode (in green). This result indicates the difficulty in disambiguating and recalling version-specific library usage. It is worth noting that the larger and latest models, such as GPT-4o (M13) and LLaMA3-70B (M12), achieve significantly better performance than the other models (See Appendix D.1 for the error analysis of GPT-4o.). However, the performance gap with HumanEval and MBPP is still large, with at least 15 points. Thus, for the simplest token-level completion task, state-of-the-art LLMs struggle to achieve satisfactory performance.
Differences in LLM performance across different data sources. We present the Pass@1 results, categorized by data sources, of token-level code completion for LLMs on VersiCode in Figure 4-a2. Comparing these three data sources, most models perform much better on Stack Overflow than on the other two, especially the source code from downstream applications. This result may be due to the high diversity in downstream applications, which requires a strong ability to tackle them. It may also suggest that Stack Overflow is highly represented in the pre-training data of LLMs, hence a greater chance of data leakage. Similar to Figure 4-a1, the outliers are still GPT-4o (M13) and LLaMA3-70B (M12), which excel in dealing with downstream applications increasing the likelihood of models memorizing specific content. Please refer to Appendix E.1 for full numeric results.
Challenges in casual intermediate library versions. We present the token-level Pass@1 results categorized by lifespan features: addition (in blue), deprecation (in orange), and general (referring to intermediate versions; in green) for the token-level code completion task (see Figure 4-b). Most models perform well in the cases of addition and deprecation due to newly added or deprecated APIs, as these versions are likely emphasized in documentation or by the community. However, most models struggle with reasoning and adapting to intermediate versions. When viewed alongside Figure 4-a2, it is evident that models such as LLaMA3-70B perform better in downstream applications and are also good at intermediate versions, benefiting from the diversity of use cases.
Table 3: Evaluation of multi-granularity version-specific code completion, and version-adapted code editing in both “old $\mapsto$ new” and “new $\mapsto$ old” scenarios.
Model Code Completion Code Editing (old $\mapsto$ new) Code Editing (new $\mapsto$ old)
Token-level Line-level Block-level Block-level Block-level
Pass@1 Pass@3 Pass@10 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 15.71 21.72 27.84 35.59 22.53 0 0 45.41 36.44 40.25 28.89
CodeLlama-13B-Instruct [37] 13.5 20.29 26.87 29.22 16.69 4.34 3.12 39.17 27.94 39.12 26.17
StarCoder2-15B [28] 14.55 21.35 27.59 20.37 12.46 1.36 0.79 35.59 26.72 41.41 27.64
CodeGemma-7B [7] 13.28 18.79 24.09 5.2 2.57 0.37 0.22 9.16 4.12 9.72 5.28
GPT-3.5-Turbo [34] 37.59 44.62 51.17 49.19 32.85 40.28 27.57 45.96 36.8 46.96 35.06
Llama-3-70B-Chat [33] 47.55 52.42 55.77 49.62 33.14 57.68 41.47 33.37 23.51 42.94 29.36
GPT-4o [36] 70.44 74.27 77.27 69.44 58.95 64.73 50.48 54.72 45.04 55.36 52.33
Reduced context increases error risk in code generation. Based on the token-level code completion performance of each model, we selected the top models for further analysis. We present the multi-granularity comparison in Table 3. When comparing the performance between line-level and block-level code completion, we can observe that smaller models tend to fail more frequently at generating correct code in block-level completion, due to less code context and the requirement to generate more content, which aligns with our intuition. Note that the results shown here have been filtered by grammar verification, a post-generation validation step that only counts code that successfully compiles in Python. If we remove grammar verification, the overall performance of block-level completion in Table 11 (Appendix E.3) is comparable to line-level completion in Table 3. This suggests that while the models can predict code-style content, they cannot guarantee correct programming grammar.
The context code in another version is still helpful, but its benefits are limited. The comparison between block-level code completion and block-level code editing is shown in Table 3. There is a significant improvement across most models, except for LLaMA3-70B and GPT4-o. When provided with code in another version as context (i.e. in the code editing task), these models can generate correct code with a much higher success rate. However, a bottleneck is evident in LLaMA3-70B and GPT4-o, where the code context hinders their performance compared to code completion.
Major version matters in version-aware code editing. As shown in Table 3, “old to new” editing and “new to old” demonstrate similar performance among models. As shown in Figure 4, we categorize editing instances according to their source and target versions, distinguishing between major and minor versions. It’s evident that when the major version serves as the source, the model’s editing performance is inferior compared to other scenarios.
Table 4: Performance of the model’s editing capabilities, influenced by transitions between source and target versions, such as shifts from major to minor versions or vice versa.
[…]
ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 12.83 11.10 23.73 18.91 30.71 22.97 36.29 21.82
CodeLlama-13B-Instruct [37] 15.4 14.2 23.07 18.29 29.10 20.65 34.09 20.36
StarCoder2-15B [28] 14.94 13.40 19.90 14.87 24.74 16.99 35.10 21.91
CodeGemma-7B [7] 7.71 5.89 7.10 6.03 6.00 3.54 13.45 7.93
GPT-3.5-Turbo [34] 14.47 12.40 24.57 20.60 32.07 25.20 39.57 25.57
Llama-3-70B-Chat [33] 15.50 14.10 20.47 17.00 21.46 16.35 32.40 19.77
GPT-4o [36] 26.21 24.60 41.39 37.23 50.01 44.20 59.55 47.40
Figure 5: Pass@1 performance for token-level code completion, grouped by year (2015-2023), with a histogram of data distribution for each year.
The programming knowledge of LLMs, particularly regarding version-specific information, is surprisingly outdated. In Figure 5 we present the Pass@1 performance for token-level code completion, grouped by year, covering 2015-2023. Additionally, we show the histogram of data distribution for each year. To ensure precise timestamps and minimize noise, we only used instances collected from library source code. As shown in Figure 5-a we can observe a general trend: the later the time, the worse the models’ performance. This is counter-intuitive compared to temporal knowledge question answering [52], where performance initially increases before declining. We further filtered for “deprecation” (Figure 5-b) and “addition” (Figure 5-c) to identify version-sensitive cases. While the sparsity of data decreases confidence in results, in both cases, we can observe a consistent decreasing trend over time. This suggests that LLMs have outdated programming knowledge, highlighting the need for rapid adaptation to newer libraries and APIs.
4 Discussion
Can we address version-controllable code generation with retrieval-augmented generation (RAG)? RAG approaches typically involve two crucial components: retrieval and in-context generation [11]. From the retrieval perspective: (1) The difficulty in disambiguating version-related queries, as embeddings for versions like "torch 2.1.3" and "torch 1.3.2" can be very similar [39]. This similarity makes it hard for models to differentiate between specific features and capabilities associated with each version. ( 2) Version information of code snippets is rarely explicitly mentioned within the code itself and may instead appear in separate configuration files like requirements.txt. This separation necessitates a more sophisticated retrieval approach, where the model must integrate information from multiple sources to accurately understand the version dependencies. However, from the perspective of in-context generation: Table 3 shows that even non-matching version contexts (i.e., code editing) can help smaller models generate grammatically correct code. The observation suggests potential for dedicated RAG approaches [18], though the benefits are limited and retrieval noise may reduce effectiveness.
How can we enhance pre-training for new code-LLMs? Figure 5 demonstrates a notable decline in the performance of all models over time. This deterioration is likely attributable to two primary factors: (1) the use of outdated pre-training data, which causes older versions of code to predominate the training set, and (2) the backward compatibility of APIs, which results in a higher prevalence of use cases and examples pertaining to older versions of these APIs [22]. To mitigate this issue and improve the models’ capabilities with newer libraries, we suggest increasing the representation of new-version codebases within the training data. This adjustment aims to enhance the models’ proficiency in utilizing contemporary libraries effectively [52, 38].
How can we address the challenge of evolving libraries in LLMs? Generating block-level or repository-level code [30] requires LLMs to understand user demands and library dependencies. Addressing this challenge involves continually training the model with new libraries using continual learning techniques [18]. These techniques enable the model to adapt to changing libraries without forgetting previously learned information. Examples include memory-based methods and various continual learning strategies [44, 46, 43]. Additionally, developing benchmark datasets that are continuously and automatically curated and maintained is crucial for evaluating the performance of models with new libraries [16]. Enriching the taxonomy [19] and maintaining datasets for evolving libraries [22] is also vital [19]. Multi-agent systems can be employed for this purpose. Aligning development and evaluation efforts will enhance the ability of LLMs in code understanding and generation capabilities, ensuring they remain effective as libraries evolve.
5 Related Work
Code Generation Models: Recent advancements in code language models [13, 7, 5, 37, 40], driven by sophisticated NLP techniques [18] and extensive code repositories [15], have resulted in substantial breakthroughs. Transformer-based large language models [32, 37, 13, 28, 5, 12, 24] have demonstrated exceptional capabilities in generating syntactically correct and semantically meaningful code from natural language descriptions. Additionally, research efforts that integrate multi-modal data [34, 36, 33], including both code and accompanying documentation [15], have significantly improved model accuracy. While in real-world software engineering,
Code Generation Datasets: The code generation [18, 40, 31] includes tasks for both code completion and code editing, ensuring comprehensive coverage of programming scenarios. Code completion [48, 49, 8, 6, 4, 14, 29, 25, 9, 27, 21, 50, 10, 53] is the task of predicting subsequent code tokens based on the given context, benefits from datasets, which provide extensive code repositories from various programming languages. These datasets enable models to learn syntactic and semantic patterns [19]. Code editing [20, 26, 55, 54, 15, 47, 2, 19, 51, 41] involves automatically generating changes to existing code, such as bug fixes or refactoring. Datasets like EvalGPTFix [51] and DebugBench [41], which focus on bug fixing and code refinement tasks, are instrumental in this area. To our knowledge, given the necessity and challenges in library evolution [18], refer to the detailed comparison in Table 1 and Appendix B.2, the proposed dataset VersiCode is the first large-scale code generation dataset, covering both code completion and code editing.
6 Conclusion
In this paper, we introduce the VersiCode dataset, a large-scale dataset of version-controllable Python code. VersiCode encompasses 300 libraries across more than 2,000 versions spanning 9 years, with a total of more than 100,000 data points. In conjunction with VersiCode, we designed two novel tasks aimed to test the capabilities of large language models in handling version-controlled code completion and editing tasks. The comprehensive experiments reveal that current models struggle significantly with these complex tasks. While token-level tasks were less challenging, block-level tasks highlighted substantial weaknesses in handling detailed coding structures. These findings highlight the need for ongoing research to better manage the dynamic nature of source code and its evolution. We plan to expand the VersiCode dataset to include more programming languages and detailed version histories. We believe version-controllable code generation is crucial and practical, and the VersiCode dataset is valuable for advancing LLMs toward deployment in professional software development.
7 Limitation
While the construction of VersiCode involved LLM-generated annotations and a partially automated pipeline, it still requires a certain amount of human annotation work for quality control. Fully automating the dataset curation and maintenance process [45] requires further development. This capability is especially important given the potential risk of dataset contamination [23] and the ever-evolving nature of libraries and APIs [22]. Another limitation of the current VersiCode dataset is its handling of multiple programming languages. While we have conducted annotations on Java, JavaScript, and C# from Stack Overflow, as shown in Appendix B.1 and  E.2, the data size for these languages is not comparable to Python. More effort, similar to that invested in constructing the Python version of VersiCode, is required for other languages. The final limitation of the current VersiCode dataset is the simplicity of the editing scheme [19], improving which may bridge the gap between code generation and real-world software engineering tasks.
Appendix A Annotation Rules
As depicted in Table 5, we have detailed the annotation and filtering rules employed during the preprocessing of raw data.
Table 5: Annotation and filtering rules.
Procedure Rules
Ranked Libraries ->StackOverflow Q&A
Filter out answers that involve the use of libraries from the ranked libraries,
and ensure these answers include content in the library version format (e.g., pandas==1.3.5) as well as code snippets.
Ranked Libraries ->Library Source Code Based on the ranked libraries, parse the source code of these libraries to find functions related to version changes.
Ranked Libraries ->Downstream Application
(1)Exclude files that do not utilize libraries and version information explicitly listed in requirements.txt.
(2)Exclude files with an average line length exceeding 100 characters.
(3)Exclude files with a maximum line length exceeding 1000 characters.
(4)Exclude files with less than 25% of alphabetic characters.
(5)Exclude files with syntax errors.
Annotation ->Metadata
StackOverflow: Filter out data that has been annotated by experts with correct library version and code snippet,
and utilize GPT to generate functionality descriptions for the code snippets.
Library Source Code: Utilize GPT to extract examples from version change function docstrings,
filter out successfully extracted data, and employ GPT to generate functionality descriptions for the examples.
Downstream Application: Utilize GPT to generate functionality descriptions for code snippets.
Appendix B Related Dataset
B.1 Detailed Data Statistics of VersiCode
As shown in Table 6, we have outlined the complete version of VersiCode in the table, which furnishes human-labeled data for three additional languages: C#, Java, and JavaScript.
Table 6: Data statistics of VersiCode, including multiple languages.
# Language Python Java C# JavaScript
# Data Source StackOverflow; Library Source Code; Downstream Application StackOverflow StackOverflow StackOverflow
# Num. of Libraries 300 19 16 33
# Num. of Versions 2,207 25 16 60
# Size of Meta Data 11,268 29 16 62
# Task Type Completion Editing (old to new) Editing (new to old) Completion Completion Completion
# Granularity Token Line Block Block Block Block Block Block
# Avg. Input Token 1,233 1,213 44 115 116 47 51 56
# Avg. Output Token 1 9 77 70 69 220 148 131
# Num. of Instances 13,532 13,530 1,618 49,346 49,346 32 21 82
B.2 Code Completion Datasets
As shown in Table 7, we compare the VersiCode-completion dataset with existing benchmarks. Note that, VersiCode stands out in annotated data size, marking it as the inaugural dataset tailored for version-specific generation.
Table 7: Comparison of VersiCode and other code completion datasets. VersiCode is the largest annotated dataset, covering multiple languages and granularities, and involving both human and LLM joint annotations.
Benchmark Source Language Samples Completion Task Granularity Collection Time Annotation
StaQC [48] StackOverflow Python, SQL 267,056 Function Programming Line-Level, Block-Level 2018 None
CoNaLa [49] StackOverflow Python, Java 2,879 Function Programming Line-Level, Block-Level 2018 Human
CT-maxmin [8] Existing Benchmark Multi(=6) 2,615 Cloze Test Token-Level 2020 None
HumanEval [6] Hand-Written Python 164 Function Programming Line-Level, Block-Level 2021 Human
MBPP [4] Hand-Written Python 974 Function Programming Block-Level 2021 Human
APPS [14] Programming Sites Python 10,000 Function Programming Line-Level, Block-Level 2021 None
CT-all [29] Existing Benchmark Multi(=6) 176,115 Cloze Test Token-Level 2021 None
CodeContests [25] Existing Benchmark, Codeforces Multi(=3) 13,610 Function Programming Block-Level 2022 None
HumanEval-FIM [9] Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2022 None
HumanEval+ [27] Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2023 Human
MBPP+ [27] Existing Benchmark Python 378 Function Programming Block-Level 2023 Human
DS-1000 [21] StackOverflow Python 1,000 Function Programming Line-Level, Block-Level 2023 Human
CoderEval [50] Github Python, Java 460 Function Programming Block-Level 2023 Human
CodeApex [10] Programming Sites C++ 476 Function Programming Block-Level 2023 None
HumanEval-X [53] Existing Benchmark Multi(=5) 820 Function Programming Line-Level, Block-Level 2023 Human
VersiCode StackOverflow, Github Python, Java, C#, JavaScript 28,682 Cloze Test, Function Programming Token-Level, Line-Level, Block-Level 2024 Human, LLM
Appendix C Evaluation Details
C.1 Hyper-parameter
As illustrated in Table 8, we have itemized the hyper-parameters pertinent to version-controllable code generation.
Table 8: Hyper-parameters for completion and editing.
hyper-parameter code completion code editing
token-level line-level block-level block-level
temperature 0.8 0.8 0.8 0.8
top_p 0.95 0.95 0.95 0.95
max_tokens 64 128 512 512
n 100 6 6 6
C.2 Prompt Template
We introduce the prompt template for token-level, line-level, and block-level evaluations in Figure 6, Figure 7, and Figure 8, respectively.
Figure 6: Prompt template for token-level version-specific code completion.
Figure 7: Prompt template for line-level version-specific code completion.
Figure 8: Prompt template for block-level version-specific code completion.
Figure 9: Prompt template for version-aware code editing.
C.3 Data Sampling
For token and line-level tasks, we randomly sampled 2,000 instances for evaluation. For block-level tasks, we used the entire dataset due to its smaller size. In the time trend experiment, we sampled 200 data points per quarter or used all available data if fewer.
Appendix D Example Analysis
D.1 Error Analysis of GPT4-o
Despite GPT4-o achieving superior performance in general evaluation, it still encounters errors in 30% of instances. We provide several negative examples in Figure 10, Figure 11, Figure 12, and Figure 13.
Figure 10: The first negative example of GPT-4o on token-level code completion.
Figure 11: The second negative example of GPT-4o on token-level code completion.
Figure 12: The third negative example of GPT-4o on token-level code completion.
Figure 13: The fourth negative example of GPT-4o on token-level code completion.
Appendix E External Experiments
E.1 A Wider Comparison Study over LLMs
In addition to the model depicted in Figure 4, comprehensive and detailed evaluation results are presented in Table 9, encompassing 23 models and sorted by the release time of each model.
Table 9: Full evaluation results of Pass@1 on token-level code completion compared to related datasets and different data sources. The results for related datasets are collected from the online leaderboard of Evalplus [27].
Release Time Model HumanEval HumanEval+ MBPP MBPP+ VersiCode
Pass@1 Pass@1 Pass@1 Pass@1 Library Source Code Downstream Application StackOverflow Total
2023.06.14 WizardCoder-15B-V1.0 [32] 56.7 50.6 64.3 54.2 0.17 0 0.1 0.06
2023.06.14 WizardCoder-Python-7B-V1.0 [32] 50.6 45.1 58.5 49.5 6.62 0.17 5.45 2.66
2023.07.18 Llama-2-7B [42] 12.8 - 20.8 - 6.57 0.46 4.76 2.74
2023.07.18 Llama-2-13B-Chat [42] 18.3 - 30.6 - 3.71 0.06 3.41 1.51
2023.08.25 CodeLlama-7B-Instruct [37] 34.8 - 44.4 - 17.77 0.62 17.8 7.62
2023.08.25 CodeLlama-13B-Instruct [37] 42.7 - 49.4 - 28.45 2.47 32.05 13.5
2023.08.28 CodeLlama-7B-Python [37] 38.4 - 47.6 - 3.4 0.03 2.35 1.28
2023.10.29 DeepSeek-Coder-6.7B-Instruct [13] 74.4 71.3 74.9 65.6 3.83 0.15 4.34 1.71
2023.11.11 Mistral-7B-Instruct-V0.2 [17] 42.1 36 44.7 37 13.96 1.85 20.33 7.54
2024.01.25 DeepSeek-Coder-7B-Instruct-V1.5 [13] 75.6 71.3 75.2 62.2 26.7 4.51 44.77 15.71
2024.01.25 GPT-3.5-Turbo [34] 76.8 70.7 82.5 69.7 40.55 30.48 65.95 37.59
2024.02.27 StarCoder2-7B [28] 35.4 29.9 55.4 45.6 12.21 0.32 13.02 5.27
2024.02.27 StarCoder2-15B [28] 46.3 37.8 66.2 53.1 29.7 2.9 35.79 14.55
2024.04.09 CodeGemma-7B-Instruct [7] 60.4 51.8 70.4 56.9 31.8 0.76 31.29 13.36
2024.04.09 CodeGemma-7B [7] 44.5 41.5 65.1 52.4 29.61 1.12 34.01 13.28
2024.04.10 aiXCoder-7B [3] 54.9 - 66 - 17.51 1.09 26.3 8.83
2024.04.15 aiXCoder-7B-Base [3] 43.2 - 62.2 - 20.41 0.94 26.37 9.59
2024.04.15 CodeQwen1.5-7B [5] 51.8 45.7 73.5 60.8 11.61 0.12 7.58 4.33
2024.04.15 CodeQwen1.5-7B-Chat [5] 83.5 78.7 79.4 69 12.16 0.33 9.2 4.81
2024.04.18 Llama-3-8B [33] 35.5 29.3 61.4 51.6 17.18 0.24 20.69 7.57
2024.04.18 Llama-3-8B-Instruct [33] 61.6 56.7 70.1 59.3 20.79 3.67 34.08 12.23
2024.04.18 Llama-3-70B-Chat [33] 77.4 72 82.3 69 33.76 50.93 64.35 47.55
2024.05.13 GPT-4o [36] 85.4 81.7 85.7 73.3 58.37 72.98 87.21 70.44
E.2 Multi-language Analysis
Table 10: Multi-language performance on VersiCode
Model Python Java C# JavaScript
ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 40.03 27.35 61.55 46.62 71.43 49.68 75.22 54.24
CodeLlama-13B-Instruct [37] 48.83 34.63 70.92 58.87 47.62 35.54 52.87 34.11
StarCoder2-15B [28] 39.71 27.36 38.63 27.43 33.33 28.63 60.67 39.33
CodeGemma-7B [7] 8.67 5.00 34.38 23.53 0 0 16.82 10.53
GPT-3.5-Turbo [34] 40.77 28.06 50.00 39.34 28.57 26.87 24.39 15.85
GPT-4o [36] 64.72 50.48 70.83 64.04 71.43 63.26 77.74 70.24
Llama-3-70B-Chat [33] 57.68 41.47 61.55 58.57 66.67 56.35 75.61 67.61
As depicted in Table 10, we perform the primary multi-language experiments. Counter-intuitively, the performance of LLMs in Java, JavaScript, and C# surpasses that in Python. This anomaly might be attributed to potential data leakage from the Stack Overflow dataset.
E.3 Block-level Generation without Grammar Verification
Figure 14: Python script for grammar verification.
As shown in Figure 14, we employ a straightforward yet effective function for basic grammar verification of generated Python code. Upon comparing Table 11 to Table 3, it becomes evident that the model tasked with editing, alongside reference code snippets from other versions, finds it easier to produce grammar-verified code.
Table 11: Block-level code completion and editing without grammar verification results.
Model Code Completion Code Editing (old to new) Code Editing (new to old)
Block-level Block-level Block-level
Score@6(ISM) Score@6(PM) Score@6(ISM) Score@6(PM) Score@6(ISM) Score@6(PM)
DeepSeek-Coder-7B-Instruct-V1.5 [13] 40.03 27.35 46.17 37.20 42.94 33.66
CodeLlama-13B-Instruct [37] 48.83 34.63 41.74 32.37 41.41 30.01
StarCoder2-15B [28] 39.71 27.36 40.94 30.73 44.46 31.88
CodeGemma-7B [7] 8.67 5.00 24.54 17.46 22.61 12.08
GPT-3.5-Turbo [34] 40.77 28.06 45.96 36.80 46.96 35.76
Llama-3-70B-Chat [33] 58.08 41.78 33.37 23.51 42.94 29.36
GPT-4o [36] 64.72 50.48 55.48 45.80 55.36 52.33